What is the kinetic energy of a 300g mass moving at 150,000 km/s?
1
Expert's answer
2011-06-24T07:03:23-0400
The kinetic energy of 300g mas can be calculated by the formula:
E = mv2/2 = 0.3[kg]*(1.5 x 105 [km/s] )2/2 = 6.75 x 109 Joules.